OpenStudy (anonymous):

20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

20 + 25 + 50(x - 2) = C, and C = 145 So: 20 + 25 + 50(x - 2) = 145 45 + 50x - 100 = 145 50x = 200 x = 4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome!